OpenStudy (anonymous):

Tommy packed 8 coolers. Each cooler has 14 bottles of water and 12 bottles of juice. Which equation could be used to calculate the total number of bottles in the coolers? A. 8 × 14 × 12 = b B. 8 × (14 + 12) = b C. (8 + 14) × 12 = b D. (8 + 14) × (8 + 12) = b

OpenStudy (anonymous):

medal + fan

OpenStudy (sleepyjess):

If we have 1 cooler, then there is 14+12 bottles, so how much would there be in 8 coolers?

OpenStudy (anonymous):

nothing like a bit of editing before posting a question, is there?

OpenStudy (sleepyjess):

LOL

OpenStudy (anonymous):

each cooler has \(14+12=26\) bottles and there are 8 coolers so \(8\times 26\) i guess you are supposed to say \[8\times (14+12)\] but that is silly since it is easy enough to compute